OpenStudy (anonymous):

Suppose you go shopping for a new futon bed for your apartment/home. The model you really like happens to be on sale for $200. It's original price is $500. What percent of the original price will you save if you purchase it? Answer is in decimal. I do not understand this please help?!!!!!

jimthompson5910 (jim_thompson5910):

You save 500 - 200 = 300 dollars The savings of the original price is 300/500 = 3/5 = 0.6, which is 60% savings The answer will be given as 0.6 since they want a decimal number

OpenStudy (anonymous):

THANK YOU THANK YOU THANK YOUUUUUUUUUUUUUUUUUUUUUUUUUUUUUUUUUU